OpenStudy (jack1):

differential equation help? how do i rearrange this eqn so it can fit the standard form of y' + P(x) y = q(x) i cant seem to isolate the y's

OpenStudy (jack1):

(1-2x^2-2y) dy/dx = (4x^3 - 4xy)

OpenStudy (jack1):

trying to do : (4x^3 - 4xy) / (1-2x^2-2y) havve 4x^3 / (1-2x^2-2y) - 4xy/ (1-2x^2-2y)

OpenStudy (jack1):

too many y's

OpenStudy (cwrw238):

i can't see that this is possible perhaps another method is required

OpenStudy (jack1):

is there another way to differentiate without finding integrating factor...?

OpenStudy (anonymous):

yeah, i think you need to do Integrating factors since My = Nx

OpenStudy (jack1):

... My = Nx...? and i cant isolate y so i cant get integrating factor...

OpenStudy (anonymous):

I may be wrong, but M(x,y) + N(x,y)*y' = 0 So M = 4x^3 - 4xy My = -4x N = (1-2x^2 - 2y) Nx = -4x So there is an exact solution

OpenStudy (anonymous):

I'm a little fuzzy on this part, but I believe you do \[Q = \int\limits_{}^{} M dx \] and then \[Q = \int\limits_{}^{} N dy\] and solve for Q. But I may have some of my subscripts backwards. I can try and sort it out if this doesn't make sense.

OpenStudy (jack1):

sorry dude, u lost me at My = -4x...?

OpenStudy (anonymous):

dM / dy = My dN/ dx = Nx

OpenStudy (jack1):

ah, cool, back woth u now

OpenStudy (anonymous):

If My = Nx, then there is an exact solution

OpenStudy (jack1):

wukk try working through it that way, cheers for that @Xetion

OpenStudy (jack1):

*will try

OpenStudy (anonymous):

I got \[x^4 - 2x^2y +y - y^2 =0\]